A particle of mass M, initially at rest, decays into two particles with rest masses
m1
and
m2
respectively. Show that the total energy of the mass
m1
is:
M
c M m m
E
2
[ ]
2
2
2
1
2 2
1
 
1
Expert's answer
2020-04-08T10:29:53-0400

"E_1+E_2=m_0c^2"


"\overrightarrow{p_1}+\overrightarrow{p_2}=0"


"E_1^2=m_1^2c^4" and "p_1^2=2m_1E_1"


"(m_0c^2-E_1)^2=E_2^2"



"(m_0c^2-E_1)^2="


"=(m_0^2c^4+m_1^2c^4-2m_0c^2m_1c^2)="


"=c^4(m_0^2+m_1^2-2m_0m_1)="


"=c^4(m_0^2+m_1^2)-c^22m_0E_1"



"E_2^2=m_2^2c^4"


"m_2^2c^4=c^4(m_0^2+m_1^2)-c^22m_0E_1\to m_2^2c^2=c^2(m_0^2+m_1^2)-2m_0E_1"


So, we have


"E_1=\frac{c^2}{2m_0}(m_0^2+m_1^2-m_2^2)" Answer.